# Preference Optimization for Review Question Generation Improves Writing Quality

### v1

@@ Abstract
Peer review relies on substantive, evidence-based questions, yet existing LLM-based approaches often generate surface-level queries, drawing over 50% of their question tokens from a paper’s first page. To bridge this gap, we develop IntelliReward, a novel reward model built from a frozen autoregressive LLM with trainable multi-head transformers over the final 50 token states, which outperforms API-based SFT baselines in predicting expert-level human preferences. By applying Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) with IntelliReward, we train IntelliAsk, a question-generation model aligned with human standards of effort, evidence, and grounding. We find consistent improvements on reasoning and writing benchmarks, suggesting reviewer-question quality correlates with broader capabilities. Compared to the Qwen3-32B base model, IntelliAsk shows measurable gains across diverse benchmarks, specifically improving performance on reasoning tasks like MuSR (68.3 vs 64.7 Acc) and complex writing evaluations such as WritingBench (8.31 vs 8.07). We release our implementation, expert preference annotations, and the IntelliReward model to provide an automatic evaluation benchmark for grounding, effort, and evidence in LLM-generated review questions.(code available at https://anonymous.4open.science/r/IntelliA-3E09/).
Preference Optimization for Review Question Generation Improves Writing Quality
Karun Sharma, Vidushee Vats, Shengzhi Li, Yuxiang Wang, Zhongtian Sun, Prayag Tiwari
@@ 3 Benchmarking SOTA Reasoning LLMs Against Humans
LLMs are capable of generating reviews when provided with a complete paper, however, they tend to fall short is in asking compelling questions that involve critical thinking about the content of the paper and as well as the domain knowledge of the paper under consideration. To study this, we conduct a human annotation study comparing questions extracted from OpenReview reviews with those generated by several state-of-the-art LLMs.
@@ 3.2 Rubrics For Assessing Question Quality: Effort, Evidence, and Grounding
First Page Bias We measure the fraction of words in the question that originate from the paper’s first page. This tests whether models rely disproportionately on introductory text when framing questions. A high score indicates surface-level dependence, while lower scores suggest engagement with the full paper. Qwen2.5-32B shows the strongest dependence, with 55% of question words coming from the first page alone (Table 3). In contrast, Human-authored questions, o3, and Gemini 2.5 Pro achieve relatively low scores, indicating that they draw more evenly from later sections of the paper when constructing questions.
@@ 4 SFT on Filtered Human Questions
We fine-tuned Qwen/Qwen2.5-7B-Instruct-1M on our curated training data using filtered Human-authored questions as the reference for reviewer-style generation. Training ran on four H200 GPUs for 24 hours with an input length of 14K tokens per paper. For evaluation, we held out a test split of 2200 samples from the curated data and used the same prompts as in our human annotation study to ensure fairness.
@@ 5.2 Reward Model Architecture and Training
Frozen None 61 64 61 62
Frozen Pool50 64 67 64 65
Train None 64 65 60 63
Train Pool50 65 69 67 67
Train Pool128 64 68 66 66
@@
Frozen None 68 68 70 69
Frozen Pool50 70 76 70 72
Frozen Pool128 69 77 67 71
Train None 71 69 70 70
Train Pool50 71 78 70 73
Train Pool128 70 78 68 72
Table 2: Ablation study comparing head architectures. Base: Frozen = Frozen backbone, Train = Trainable backbone. Pool: Pooling strategy (kk=last $k$ tokens). Scores: Eff.=Effort, Evid.=Evidence, Grd.=Grounding.
@@ 6.2 Automatic Evaluation with IntelliReward
Table 4: Automatic evaluation using IntelliReward on test set. IntelliAsk-32B achieves the highest score among small models (0.55/3.0), substantially outperforming SFT-only baselines. Among all models, o3 achieves the best performance. Bold: best in category; underline: second-best. FPB = First Page Bias
To enable scalable evaluation, we use IntelliReward to automatically assess question quality on the entire test set. Table 4 reports performance across both large proprietary models and smaller open-source models. The automatic evaluation reveals several key findings:
Large Models: Among frontier models, o3 with medium reasoning achieves the strongest performance (0.72/3.0), with particularly high scores on Effort (0.28) and Grounding (0.30). GPT-5 and Gemini 2.5 Pro demonstrate moderate performance (0.45–0.51), while other models like Claude 3.7 and GPT-4.1 score lower.
Small Models: IntelliAsk-32B substantially outperforms other small models ($\leq$32B parameters), achieving 0.55/3.0 compared to 0.03–0.10 for SFT baselines (Qwen2.5-7B SFT, OpenReviewer, DeepReviewer). This gap demonstrates the value of reinforcement learning with human preference alignment over pure supervised fine-tuning. Our smaller IntelliAsk-7B variant achieves 0.17/3.0, showing promise for resource-constrained settings.
First Page Bias Analysis: Models with lower first page bias tend to produce more thoughtful questions. IntelliAsk-32B achieves 21.37% first page bias, substantially lower than SFT baselines (42–51%), indicating better engagement with the full paper content.
@@ 8 Conclusion
The dataset was created from reviewer comments on ICLR papers that are publicly available on Openreview.net. We restricted the collection to text that is already accessible to the public and removed any metadata that could identify reviewers. As OpenReview content is distributed under the Creative Commons Attribution 4.0 International (CC BY 4.0) license, our use and release of these comments complies with the original license terms. The human preference dataset was constructed through additional human annotation on top of the collected review comments. These annotations are original contributions by our paper and are released under the same CC BY 4.0 license.
We do not claim copyright over the original review texts or paper excerpts used in our datasets.
@@ A.5 Examples of Questions Generated from Openreviewer, DeepReviewer and IntelliAsk
: Concerning the computation requirement, one gradient descent step seems to become at least slightly more computational expensive as compared to DQN. In the paper it is mentioned at the very end of the discussion section that "with K=5 (it) only requires 1 to 2 times more time to run". Can this quantification be more accurate or why would it "sometimes" same compute time and "sometimes" double the compute time?
@@ A.14 Examples of Effortful, Substantive, and Evidence-Based Questions
Evidence Q: ‘This way, we transform… bypassing the time-consuming gradient computation…’ – For MINE, we do need to update NNs’ parameters. But InfoNet also needs gradient ascent. How to understand ‘bypassing the time-consuming gradient computation’?
@@ A.17.4 Extraction of Questions
By Questions I mean all the questions -- explicitly or implicitly asked -- that the author needs to answer the reviewer based on the review text.
@@
"I have questions about the learning process of the $\times$1$ conv layer in equation (5). How is it exactly trained? And is it sensitive to the training sample size?\
@@
\"Question\": ""I have questions about the learning process of the $\times$1$ conv layer in equation (5). How is it exactly trained? And is it sensitive to the training sample size?"
@@
"There should be a discussions of assumptions behind the OC-GFNs pretraining. Namely, that transfer is only possible when the reward function changes but not if the action-space or the state-space change. Moreover, the goal-conditioning requires a well specified set of outcomes Y -- presumably not all states s are terminal states -- which makes the proposed method not truly unsupervised. These limitations (together with the applicability mentioned at the end of A.2) could be stated explicitly in the main text, and left to future work.\
@@
\"Question\": "While there are enough benchmarks, I believe none include continuous action/state spaces. Moreover, the experiments only one GFN variant -- the detailed-balance one, which is also used for OC-GFN. It would help validate the generality of OC if we had experiments showing it worked on these different settings"

### v2

# IntelliAsk: Learning to Ask High-Quality Research Questions via RLVR

@@ Abstract
Peer review relies on substantive, evidence-based questions, yet current LLMs generate surface-level queries that perform worse than human reviewer questions in expert evaluation. To address this gap, we curate a high-quality dataset of reviewer questions from OpenReview and conduct a human preference study where expert annotators evaluate question-paper pairs across three dimensions: effort, evidence, and grounding. From these annotations, we train IntelliReward, a reward model built from a frozen autoregressive LLM with trainable multi-head transformers. Validated against expert judgments, IntelliReward predicts reviewer-question quality better than API-based SFT baselines and provides scalable evaluation. We apply Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) with IntelliReward to train IntelliAsk, a question-generation model aligned with human standards of effortful, evidence-based critique. Human evaluations show IntelliAsk generates more grounded, substantive and effortful questions than strong baselines and reduces reliance on first-page content. We also find improvements on reasoning and writing benchmarks, suggesting reviewer-question quality correlates with broader capabilities. Compared to Qwen3-32B, IntelliAsk improves MuSR (68.3 vs 64.7 Acc) and WritingBench (8.31 vs 8.07). We release our code, filtered review dataset, expert annotations, IntelliAsk and IntelliReward to support automatic evaluation of grounding, effort, and evidence in LLM-generated review questions. (https://anonymousse123456.github.io/intelliask.github.io/ ).
IntelliAsk: Learning to Ask High-Quality Research Questions via RLVR
Karun Sharmathanks: Equal contribution., Vidushee Vats11footnotemark: 1, Shengzhi Li11footnotemark: 1, Yuxiang Wang, Zhongtian Sun, Prayag Tiwari
@@ 3 Benchmarking SOTA Reasoning LLMs Against Humans
LLMs are capable of generating reviews when provided with a complete paper, however, they tend to fall short in asking compelling questions that involve critical thinking about the content of the paper and as well as the domain knowledge of the paper under consideration. To study this, we conduct a human annotation study comparing questions extracted from OpenReview reviews with those generated by several state-of-the-art LLMs.
@@ 3.2 Rubrics For Assessing Question Quality: Effort, Evidence, and Grounding
First Page Bias (FPB). We measure the fraction of words in the question that originate from the paper’s first page. This tests whether models rely disproportionately on introductory text when framing questions. A high score indicates surface-level dependence, while lower scores suggest engagement with the full paper. Qwen2.5-32B shows the strongest dependence, with 55% of question words coming from the first page alone (Table 3). In contrast, Human-authored questions, o3, and Gemini 2.5 Pro achieve relatively low scores, indicating that they draw more evenly from later sections of the paper when constructing questions. FPB is used only as an evaluation metric and is not part of the training objective.
@@ 4 SFT on Filtered Human Questions
We fine-tuned Qwen/Qwen2.5-7B-Instruct-1M on our curated training data using filtered Human-authored questions as the reference for reviewer-style generation. Training ran on four H200 GPUs for 24 hours with an input length of 14K tokens per paper. For evaluation, we held out a test split of 2200 samples from the curated data and used the same prompts as in our human annotation study to ensure fairness. During human evaluation, we observed that the SFT-generated questions were generally weak, we therefore additionally report automatic evaluation scores using our reward model, IntelliReward (see Section 5.1 for IntelliReward).
@@ 5.2 Reward Model Architecture and Training
[Uncaptioned image] None 61 64 61 62
[Uncaptioned image] Pool50 64 67 64 65
[Uncaptioned image] None 64 65 60 63
[Uncaptioned image] Pool50 65 69 67 67
[Uncaptioned image] Pool128 64 68 66 66
@@
[Uncaptioned image] None 68 68 70 69
[Uncaptioned image] Pool50 70 76 70 72
[Uncaptioned image] Pool128 69 77 67 71
[Uncaptioned image] None 71 69 70 70
[Uncaptioned image] Pool50 71 78 70 73
[Uncaptioned image] Pool128 70 78 68 72
Table 2: Ablation study comparing head architectures. Base: [Uncaptioned image] = Frozen backbone, [Uncaptioned image] = Trainable backbone. Pool: Pooling strategy (kk=last $k$ tokens). Scores: Eff.=Effort, Evid.=Evidence, Grd.=Grounding.
@@ 6.2 Automatic Evaluation with IntelliReward
Table 4: Automatic evaluation using IntelliReward on test set. Rows highlighted in beige correspond to SFT baseline models (OpenReviewer-8B, DeepReviewer-7B, and Qwen2.5-7B SFT). IntelliAsk-32B achieves the highest score among small models (0.55/3.0), substantially outperforming SFT-only baselines. Among all models, o3 achieves the best performance. Bold: best in category; underline: second-best. FPB = First Page Bias
@@ 8 Conclusion
The dataset was created from publicly available reviewer comments on ICLR papers hosted on Openreview.net. We restricted the collection to publicly accessible text and removed any metadata that could identify reviewers. As OpenReview content is distributed under the CC BY 4.0 license, our use and release of these comments complies with the original terms. The human preference annotations are original contributions and are released under the same CC BY 4.0 license. We do not claim copyright over the original review texts or paper excerpts.
@@ A.5 Examples of Questions Generated from Openreviewer, DeepReviewer and IntelliAsk
: Concerning the computation requirement, one gradient descent step seems to become at least slightly more computationally expensive as compared to DQN. In the paper it is mentioned at the very end of the discussion section that "with K=5 (it) only requires 1 to 2 times more time to run". Can this quantification be more accurate or why would it "sometimes" same compute time and "sometimes" double the compute time?
@@ A.14 Examples of Effortful, Substantive, and Evidence-Based Questions
Evidence Q: ‘This way, we transform… bypassing the time-consuming gradient computation…’ — For MINE, we do need to update NNs’ parameters. But InfoNet also needs gradient ascent. How to understand ‘bypassing the time-consuming gradient computation’?
@@ A.17.4 Extraction of Questions
By Questions I mean all the questions --- explicitly or implicitly asked --- that the author needs to answer the reviewer based on the review text.
@@
"I have questions about the learning process of the 1x1 conv layer in equation (5). How is it exactly trained? And is it sensitive to the training sample size?\
@@
\"Question\": ""I have questions about the learning process of the 1x1 conv layer in equation (5). How is it exactly trained? And is it sensitive to the training sample size?"
@@
"There should be a discussions of assumptions behind the OC-GFNs pretraining. Namely, that transfer is only possible when the reward function changes but not if the action-space or the state-space change. Moreover, the goal-conditioning requires a well specified set of outcomes Y --- presumably not all states s are terminal states --- which makes the proposed method not truly unsupervised. These limitations (together with the applicability mentioned at the end of A.2) could be stated explicitly in the main text, and left to future work.\
@@
\"Question\": "While there are enough benchmarks, I believe none include continuous action/state spaces. Moreover, the experiments only one GFN variant --- the detailed-balance one, which is also used for OC-GFN. It would help validate the generality of OC if we had experiments showing it worked on these different settings"